OpenStudy (anonymous):

–21 – (–13) A. –34 B. –8 C. 8 D. 34

OpenStudy (anonymous):

34?

OpenStudy (anonymous):

the two -'s right by eachother in the middle of the equation turn into a + always.

OpenStudy (anonymous):

so it would then be -21+13

OpenStudy (anonymous):

ok but what is the perenthiesais for

OpenStudy (anonymous):

mostly for organization. I choose to ignore the, the don't change your answer or anything.

OpenStudy (anonymous):

but we got to add so shouldnt it be -34

OpenStudy (anonymous):

21+13 is 34

OpenStudy (anonymous):

The negative at the very front of the equation doesn't change. Just the two in the middle.

OpenStudy (anonymous):

-34

OpenStudy (anonymous):

-21-(-13)= -21+13= -8 Plz just trust me on this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yyyaaasss

OpenStudy (anonymous):

:)